OpenStudy (anonymous):

Ok 5(x^-3)(y^0) I cant have any Zero, negative or fraction exponents.

OpenStudy (anonymous):

That just simplifies to

OpenStudy (anonymous):

Woops: Latex isn't working...\[5/(x^3)\]

OpenStudy (anonymous):

\[y^0=1\] \[x^{-3}=\frac{1}{x^3}\] \[5=5\] \[\frac{5}{x^3}\]

OpenStudy (anonymous):

^ Sexy answer

OpenStudy (anonymous):

thnx

OpenStudy (anonymous):

learning latex. you?

OpenStudy (anonymous):

Latex?

OpenStudy (anonymous):

I got latex... it just wasn't working :D

OpenStudy (anonymous):

how i write \[x^3\] etc

OpenStudy (anonymous):

Is it a program?

OpenStudy (anonymous):

i am trying to learn it. just typing it in to see if i can get faster and remember commands not using an editor although i use one for writing in general

OpenStudy (anonymous):

it is a way to typeset math but it is not a wisyisyg so not so easy to learn. not like typing in microdemon word.